OpenStudy (annonomus):

There are (36)2 ⋅ 30 candies in a store. What is the total number of candies in the store? A ) 3^3 B ) 3^4 C ) 3^8 D ) 3^12

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\) try tagging some1 to your q like this @annonomus

OpenStudy (welshfella):

is is 

OpenStudy (annonomus):

yes

OpenStudy (welshfella):

36^2 = (6^2)^2 = 6^4

OpenStudy (welshfella):

6^4 + 30 = 38880 which is not an option Are you sure your question is correct?